OpenStudy (jkbo):

The area of a triangle with base 19 in. and height 11 in. is _____. Be sure to include units.

OpenStudy (anonymous):

just use the area equation of a triangle. A = (1/2)*base*height your height is 11 base is 19. plug in! :)

OpenStudy (lgbasallote):

\[\large area = \frac{base \times height}{2}\]

OpenStudy (jkbo):

thanks